OpenStudy (anonymous):

What is the solution of the system ?Use substitution 7x-3y=2 -2y+x=-6

OpenStudy (mathstudent55):

Solve one equation for one variable. Then substitute what that variable is equal to in the other equation. In this case, the second equation is easy to solve for x: -2y + x = -6 Add 2y to both sides: x = 2y - 6 Now that we have an equartion solved for x, and x = 2y - 6, we insert 2y - 6 in for x in the first equation: 7x - 3y = 2, insert 2y - 6 in for x: 7(2y - 6) - 3y = 2 Distribute the 7: 14y - 42 - 3y = 2 Combine like terms: 11y = 44 y = 4 Now insert the value of y in either of the original equation to get x. Let's use the second equation: -2y + x = -6, but y = 4, so -2(4) + x = -6 -8 + x = -6 Add 8 to both sides: x = 2 Solution: x = 2, y = 4

OpenStudy (anonymous):

Wow, THANK YOU so much. I was so confused

OpenStudy (mathstudent55):

You WERE confused. Did this help?